the dolphin jumps out the water at a show, the function can be modeled by
f(x)=-16x^2 times 32x, which the x is the time in second. find out what is the max. height, and how long does it take to reach that height, and how long will the dolphine stay in the air before he went back into the water.

Answer by Alan3354(69443)

the dolphin jumps out the water at a show, the function can be modeled by
f(x)=-16x^2 times 32x, which the x is the time in second. find out what is the max. height, and how long does it take to reach that height, and how long will the dolphine stay in the air before he went back into the water.
----------
f(x) = -16x^2 + 32x (not times)
max height is the vertex of the parabola
at x = -b/2a = -32/(-32) = 1 second to max height
Then 1 second to fall = 2 seconds total in air.
--------------------
f(1) = -16 + 32 = 16 feet max height